OpenStudy (anonymous):

in which are all the solutions to this inequality x >5 x - 8 A. { -1, 0, 1, 2 } B. { 0, 1, 2, 3 } C. { 3, 4, 5, 6 } D. { -2, -1, 0 1 }

OpenStudy (anonymous):

first, lets isolate the constants together  now let's start plugging in numbers from the set, { -1, 0, 1, 2 }, to see if the set is true for false (x)-5(x)>-8 let x= -1 (-1)-5(-1)>-8 -1+5>-8 4>-8 [true] let x = 0 (0)-5(0)>-8 [true] let x = 1 (1)-5(1)>-8 [true] let x =2 (2)-5(2)>-8 [false] so this set of { -1, 0, 1, 2 } is not a solution set. do this same method for the rest of the choices.